# Bridging the performance-gap between target-free and target-based reinforcement learning

### v2

@@ Abstract
The use of target networks in deep reinforcement learning is a widely popular solution to mitigate the brittleness of semi-gradient approaches and stabilize learning. However, target networks notoriously require additional memory and delay the propagation of Bellman updates compared to an ideal target-free approach. In this work, we step out of the binary choice between target-free and target-based algorithms. We introduce a new method that uses a copy of the last linear layer of the online network as a target network, while sharing the remaining parameters with the up-to-date online network. This simple modification enables us to keep the target-free’s low-memory footprint while leveraging the target-based literature. We find that combining our approach with the concept of iterated $Q$-learning, which consists of learning consecutive Bellman updates in parallel, helps improve the sample-efficiency of target-free approaches. Our proposed method, iterated Shared $Q$-Learning (iS-QL), bridges the performance gap between target-free and target-based approaches across various problems, while using a single $Q$-network, thus being a step forward towards resource-efficient reinforcement learning algorithms.
@@ 1 Introduction
Originally, Q-learning (Watkins & Dayan, 1992) was introduced as a reinforcement learning (RL) method that performs asynchronous dynamic programming using a single look-up table. By storing only one $Q$-estimate, Q-learning benefits from an up-to-date estimate and a low memory footprint. However, replacing look-up tables with non-linear function approximators and allowing off-policy samples to make the method more tractable introduces training instabilities (Sutton & Barto, 2018). To address this, Mnih et al. (2015) introduce Deep Q-Network (DQN), an algorithm that constructs the regression target from an older version of the online network, known as the target network, which is periodically updated to match the online network (see “Target Based” in Figure 1). This modification to the temporal-difference objective helps mitigate the negative effects of function approximation and bootstrapping (Zhang et al., 2021), two elements of the deadly triad (van Hasselt et al., 2018). Recently, new methods have demonstrated that increasing the size of the $Q$-network can enhance the learning speed and final performance of temporal difference methods (Espeholt et al., 2018; Schwarzer et al., 2023; Nauman et al., 2024; Lee et al., 2025). Numerous ablation studies highlight the crucial role of the target network in maintaining performance improvements over smaller networks (Figure $7$ in Schwarzer et al. (2023), and Figure $9b$ in Nauman et al. (2024)). Interestingly, even methods initially introduced without a target network (Bhatt et al. (2024) and Kim et al. (2019)) benefit from its reintegration (Figure $5$ in Palenicek et al. (2025) and Gan et al. (2021)).
@@
While temporal difference methods clearly benefit from target networks, their utilization doubles the memory footprint dedicated to $Q$-networks. This ultimately limits the size of the online network due to the constrained Video Random Access Memory (VRAM) of GPUs. This limitation is not only problematic for learning on edge devices where memory is constrained, but also for applications that inherently require large network sizes, such as handling high-dimensional state spaces (Boukas et al., 2021; Pérez-Dattari et al., 2019), processing multi-modal inputs (Schneider et al., 2025), or constructing mixtures of experts (Obando Ceron et al., 2024; Hendawy et al., 2024). This motivates the development of target-free methods (see “Target Free” in Figure 1).
@@
Figure 1: We propose a simple alternative to target-based/target-free approaches, where a linear layer represents the target network, sharing the rest of the parameters with the online network (Shared Features). We apply the concept of iterated Q-learning (Vincent et al., 2025), which consists of learning multiple Bellman updates in parallel, to reduce the performance gap between target-free and target-based approaches (iterated Shared Features).
@@
In this work, we introduce an alternative to the binary choice between target-free and target-based approaches. We propose storing only the smallest possible part of the target network, i.e., the parameters of the last linear layer, while using the parameters of the online network to substitute the other layers of the target network (see “Shared Features” in Figure 1). Although this simple modification alone helps reduce the performance gap between target-free and target-based DQN (see “iS-DQN K=1K=1” in Figure 4, right), we explain in this work how it opens up the possibility of leveraging the target-based literature to reduce this gap further, while maintaining a low memory footprint. Notably, this approach is also orthogonal to regularization techniques that have been shown to be effective for target-free algorithms (Kim et al., 2019; Bhatt et al., 2024; Gallici et al., 2025). Therefore, we will build upon these approaches to benefit from their performance gains.
@@
In the following, we leverage the concept of iterated Q-learning (Vincent et al., 2025) to enhance the learning speed (in terms of number of environment interactions) of target-free algorithms. This concept, initially introduced as a target-based approach, aims at learning multiple Bellman iterations in parallel. This leads to a new algorithm, termed iterated Shared Q-Network (iS-QN), pronounced “ice-QN” to emphasize that it contains a frozen head. iS-QN utilizes a single network with multiple linear heads, where each head is trained to represent the Bellman target of the previous one (see “iterated Shared Features” in Figure 1). Our evaluation of iS-QN across various RL settings demonstrates that it improves the learning speed of target-free methods while maintaining a comparable memory footprint and training time.
@@ 2 Background
Deep $Q$-Network (Mnih et al., 2015)  The optimal policy of a Markov Decision Process (MDP) with a discrete action space can be obtained by selecting for each state, the action that maximizes the optimal action-value function $Q^{*}$. This function represents the largest achievable expected sum of discounted rewards given a state-action pair. This is why Mnih et al. (2015) approximate the optimal action-value function with a neural network $Q_{\theta}$, represented by a vector of parameters $\theta$. This neural network is learned to approximate its Bellman iteration $\Gamma Q_{\theta}$, leveraging the contraction property of the Bellman operator $\Gamma$ to guide the optimization process toward the operator’s fixed point, i.e., the optimal action-value function $Q^{*}$. In practice, a sample estimate of the Bellman iteration is used, where for a sample $(s,a,r,s^{\prime})$, $\Gamma Q_{\theta}(s,a)=r+\gamma\max_{a^{\prime}}Q_{\theta}(s^{\prime},a^{\prime})$, where $\gamma$ is the discount factor linked to the MDP of interest. However, this learning procedure is unstable because the neural network $Q_{\theta}$ learns from its own values, which change at each optimization step due to function approximation, and because of the compound effect of the overestimation bias. To mitigate these issues, the authors introduce a target network with parameters $\bar{\theta}$ to stabilize the regression target $\Gamma Q_{\bar{\theta}}$, and periodically update these parameters to the online parameters $\theta$ every $T$ steps. On the negative side, this doubles the memory footprint dedicated to $Q$-networks.
@@
Iterated Q-Network (Vincent et al., 2025)  By using a target network, DQN slows down the training process as multiple gradient steps are dedicated to each Bellman iteration, as $\Gamma Q_{\bar{\theta}}$ is delayed by some gradient steps compared to $\Gamma Q_{\theta}$. To increase the learning speed, Vincent et al. (2025) propose to learn consecutive Bellman iterations in parallel. This approach uses a sequence of online parameters $(\theta_{i})_{i=1}^{K}$ and a sequence of target parameters $(\bar{\theta}_{i})_{i=0}^{K-1}$. Each online network $Q_{\theta_{i+1}}$ is trained to regress $\Gamma Q_{\bar{\theta}_{i}}$. Similarly to DQN, each target parameter $\bar{\theta}_{i}$ is updated to the online parameter $\theta_{i+1}$ every $T$ steps. Importantly, the structure of a chain is enforced by setting each $\bar{\theta}_{i}$ to $\theta_{i}$ every $D\ll T$ steps so that each $Q_{\theta_{i+1}}$, which is learned to regress $\Gamma Q_{\bar{\theta}_{i}}$, are forced to approximate $\Gamma Q_{\theta_{i}}$. This results in $Q_{\theta_{K}}\approx\Gamma Q_{\theta_{K-1}}\approx\ldots\approx\Gamma^{K}Q_{\theta_{0}}$, thus learning $K$ consecutive Bellman iterations in parallel. Importantly, DQN can be recovered by setting $K=1$. While the feature representation can be shared across the online $Q$-networks, iterated $Q$-Network (i-QN) has the drawback of requiring an old copy of the online networks to stabilize training, significantly increasing the memory footprint. In the following, we will explain how the concept of i-QN can help reduce the performance gap between target-free and target-based approaches while maintaining a low memory footprint.
@@ 3 Related Work
Alternatively, some works construct the regression target from the online network instead of the target network, but still use a target network in some other way. For example, Ohnishi et al. (2019) compute the TD(0) loss from the online network and add a term in the loss to constrain the predictions of the online network for the next state-action pair $(s^{\prime},a^{\prime})$ to remain close to the one predicted by the target network. Piché et al. (2021; 2023) develop a similar approach, enforcing similar values for the state-action pair $(s,a)$. Lindström et al. (2025) show that the target network can be removed after a pretraining phase in which they rely on expert demonstrations.
@@
Many regularization techniques have been developed, attempting to combat the performance drop that occurs when removing the target network. We stress that our approach is orthogonal to these regularization techniques and we show in Section 5 that our method improves the performance of target-free methods equipped with these advancements. Li & Pathak (2021) encode the input of the $Q$-network with learned Fourier features. While this approach seems promising, the authors acknowledge that the performance degrades for high-dimensional problems. Shao et al. (2022) remove the target-network and search for an action that maximizes the $Q$-network predictions more than the action proposed by the policy. Searching for a better action requires additional resources and is only relevant for actor-critic algorithms. Kim et al. (2019) leverage the MellowMax operator to get rid of the target network. However, the temperature parameter needs to be tuned (Kim, 2020), which increases the compute budget, and a follow-up work demonstrates that the reintegration of the target network is beneficial (Gan et al., 2021). Finally, Bhatt et al. (2024) point out the importance of using batch normalization (Ioffe & Szegedy, 2015) to address the distribution shift of the input given to the critic. Our investigation reveals that it degrades the performance in a discrete action setting (see Figure 15, right).
@@
The idea of learning multiple Bellman iterations has been introduced by Schmitt et al. (2022). They demonstrate convergence guarantees in the case of linear function approximation. Then, Vincent et al. (2024) used this approach to learn a recurrent hypernetwork generating a sequence of $Q$-functions where each $Q$-function approximates the Bellman iteration of the previous $Q$-function. Finally, Vincent et al. (2025) introduced iterated $Q$-Network as a far-sighted version of DQN that learns the $K$ following Bellman iterations in parallel instead of only learning the following one. While promising, those approaches rely on a separate copy of the learnable parameters to stabilise the training process, which increases the memory footprint. In this work, we propose to leverage the potential of iterated $Q$-learning to boost the learning speed of target-free algorithms.
@@ 4 Method
Our goal is to design a new algorithm that improves the learning speed of target-free value-based RL methods without significantly increasing the number of parameters used by the $Q$-networks. To achieve this, we consider a single $Q$-network parameterized with $K+1$ heads. We note $\omega_{k}$ the parameters of the $k^{\text{th}}$ head, $\omega$ the shared parameters, and define $\theta=(\omega,\omega_{0},..,\omega_{K})$ and $\theta_{k}=(\omega,\omega_{k})$. Following Vincent et al. (2025), for a sample $d=(s,a,r,s^{\prime})$, the training loss is
@@
In the following, we apply iterated Shared Features to several target-based approaches on multiple RL settings, demonstrating that it reduces the gap between target-free and target-based methods. For each algorithm A, we note TB-A as its target-based version, TF-A as its target-free version, and iS-A as the iterated shared approach, where “iS” stands for iterated Shared. Importantly, we incorporate the insights provided by Gallici et al. (2025) to use LayerNorm (Ba et al., 2016) for the experiments with discrete action spaces, as we found it beneficial, even for the target-based approach. Similarly, we use BatchNorm (Ioffe & Szegedy, 2015), as suggested by Bhatt et al. (2024), to improve sample-efficiency in continuous action settings, except for the target-based approach, as it degrades performances (see Figure 19, right).
@@ 5.3 Online Continuous Control
We investigate the behavior of iS-QN on the DeepMind Control suite (Tassa et al., 2018), focusing on the hard tasks. We select Soft Actor-Critic (SAC, Haarnoja et al. (2018)) as the base algorithm and adapt the architecture to the one proposed by Lee et al. (2025) (SimbaV2) so that the target-based approach corresponds to the state-of-the-art. This experiment allows us to test iS-QN on different learning dynamics, as the target updates are done with an exponentially moving average instead of a hard update, and the loss for the critic uses a categorical distribution to learn the distribution of the return. Interestingly, Figure 6 (left) shows that only using an old copy of the last layer of the critic to construct the regression target (iS-SAC $K=1$) recovers the performance drop incurred by the target-free approach compared to the target-based approach. Importantly, Lee et al. (2025) design the critic with significantly more parameters than the actor, as commonly done in the actor-critic literature  (Mysore et al., 2021; Mastikhina et al., 2025). This means that iS-SAC K=1 reduces the total number of parameters by $49\%$, see Figure 6 (right). When considering more heads to learn the following Bellman updates, we find it beneficial to give more importance to the first Bellman updates by scaling the future terms in the loss by a discounting factor of $0.25$. We note that in this setting, iS-SAC $K=9$ only performs on par with $K=1$. Nonetheless, iS-SAC $K=9$ is still performing better than the best target-free approach, having overlapping confidence intervals with the target-based approach, which serves as the goal standard, as it requires additional parameters.

### v3

# arXiv:2506.04398v3[cs.LG] 27 Feb 2026

@@ Abstract
The use of target networks in deep reinforcement learning is a widely popular solution to mitigate the brittleness of semi-gradient approaches and stabilize learning. However, target networks notoriously require additional memory and delay the propagation of Bellman updates compared to an ideal target-free approach. In this work, we step out of the binary choice between target-free and target-based algorithms. We introduce a new method that uses a copy of the last linear layer of the online network as a target network, while sharing the remaining parameters with the up-to-date online network. This simple modification enables us to keep the target-free’s low-memory footprint while leveraging the target-based literature. We find that combining our approach with the concept of iterated $Q$-learning, which consists of learning consecutive Bellman updates in parallel, helps improve the sample-efficiency of target-free approaches. Our proposed method, iterated Shared $Q$-Learning (iS-QL), bridges the performance gap between target-free and target-based approaches across various problems while using a single $Q$-network, thus stepping towards resource-efficient reinforcement learning algorithms.
@@ 1 Introduction
Originally, $Q$-learning (Watkins and Dayan, 1992) was introduced as a reinforcement learning (RL) method that performs asynchronous dynamic programming using a single look-up table. By storing only one $Q$-estimate, $Q$-learning benefits from an up-to-date estimate and a low memory footprint. However, replacing look-up tables with non-linear function approximators and allowing off-policy samples to make the method more scalable introduces training instabilities (Sutton and Barto, 2018). To address this, Mnih et al. (2015) introduce Deep $Q$-Network (DQN), an algorithm that constructs the regression target from an older version of the online network, known as the target network, which is periodically updated to match the online network (see “Target Based” in Figure 1). This modification to the temporal-difference objective helps mitigate the negative effects of function approximation and bootstrapping (Zhang et al., 2021), two elements of the deadly triad (van Hasselt et al., 2018). Recently, new methods have demonstrated that increasing the size of the $Q$-network can enhance the learning speed and final performance of temporal difference methods (Espeholt et al., 2018; Schwarzer et al., 2023; Nauman et al., 2024; Lee et al., 2025). Numerous ablation studies highlight the crucial role of the target network in maintaining performance improvements over smaller networks (Figure $7$ in Schwarzer et al. (2023), and Figure $9b$ in Nauman et al. (2024)). Interestingly, even methods initially introduced without a target network (Bhatt et al. (2024) and Kim et al. (2019)) benefit from its reintegration (Figure $5$ in Palenicek et al. (2025) and Gan et al. (2021)).
@@
While temporal difference methods clearly benefit from target networks, their utilization doubles the memory footprint dedicated to $Q$-networks. This ultimately limits the size of the online network due to the constrained memory capacity of the processing units (e.g. the vRAM of a GPU). This limitation is not only problematic for learning on edge devices where memory is constrained, but also for applications that inherently require large network sizes, such as handling high-dimensional state spaces (Boukas et al., 2021; Pérez-Dattari et al., 2019), processing multi-modal inputs (Schneider et al., 2026), or constructing mixtures of experts (Obando Ceron et al., 2024; Hendawy et al., 2024). This motivates the development of target-free methods (see “Target Free” in Figure 1).
@@
Figure 1: We propose a simple alternative to target-based/target-free approaches, where a linear layer represents the target network, sharing the rest of the parameters with the online network (Shared Features). We apply the concept of iterated $Q$-learning (Vincent et al., 2025b), which consists of learning multiple Bellman updates in parallel, to reduce the performance gap between target-free and target-based approaches (iterated Shared Features).
@@
In this work, we introduce an alternative to the binary choice between target-free and target-based approaches. We propose storing only the parameters of the last linear layer, while using the parameters of the online network to substitute the other layers of the target network (see “Shared Features” in Figure 1). Although this simple modification alone helps reduce the performance gap between target-free and target-based DQN (see “iS-DQN K=1K=1” in Figure 4, right), we explain in this work how it opens up the possibility of leveraging the target-based literature to reduce this gap further, while maintaining a low memory footprint. Notably, this approach is also orthogonal to regularization techniques that have been shown to be effective for target-free algorithms (Kim et al., 2019; Bhatt et al., 2024; Gallici et al., 2025). Therefore, we will build upon these approaches to benefit from their performance gains.
@@
In the following, we leverage the concept of iterated $Q$-learning (Vincent et al., 2025b) to enhance the learning speed (in terms of the number of environment interactions) of target-free algorithms, which is a major bottleneck in many real-world applications. The concept of iterated $Q$-learning, initially introduced as a target-based approach, aims at learning multiple Bellman iterations in parallel. This leads to a new algorithm, termed iterated Shared $Q$-Network (iS-QN), pronounced “ice-QN” to emphasize that it contains a frozen head. iS-QN utilizes a single network with multiple linear heads, where each head is trained to represent the Bellman target of the previous one (see “iterated Shared Features” in Figure 1). Our evaluation of iS-QN across various RL settings demonstrates that it improves the learning speed of target-free methods while maintaining a comparable memory footprint and training time.
@@ 2 Background
Deep $Q$-Network (Mnih et al., 2015)  The optimal policy of a Markov Decision Process (MDP) can be obtained by selecting for each state, the action that maximizes the optimal action-value function $Q^{*}$. This function represents the largest achievable expected sum of discounted rewards given a state-action pair. In the context of discrete action spaces, Mnih et al. (2015) approximate the optimal action-value function with a neural network $Q_{\theta}$, represented by a vector of parameters $\theta$. This neural network is learned to approximate its Bellman iteration $\Gamma Q_{\theta}$, leveraging the contraction property of the Bellman operator $\Gamma$ in the value space to guide the optimization process toward the operator’s fixed point, i.e., the optimal action-value function $Q^{*}$. In practice, a sample estimate of the Bellman iteration is used, where for a sample $(s,a,r,s^{\prime})$, $\Gamma Q_{\theta}(s,a)=r+\gamma\max_{a^{\prime}}Q_{\theta}(s^{\prime},a^{\prime})$, where $\gamma$ is the discount factor. However, this learning procedure is unstable because the neural network $Q_{\theta}$ learns from its own values, which change at each optimization step due to function approximation, and because of the compound effect of the overestimation bias. To mitigate these issues, the authors introduce a target network with parameters $\bar{\theta}$ to stabilize the regression target $\Gamma Q_{\bar{\theta}}$, and periodically update these parameters to the online parameters $\theta$ every $T$ steps. On the negative side, this doubles the memory footprint dedicated to $Q$-networks.
@@
Iterated $Q$-Network (Vincent et al., 2025b)  By using a target network, DQN slows down the training process as multiple gradient steps are dedicated to each Bellman iteration, as $\Gamma Q_{\bar{\theta}}$ is delayed by some gradient steps compared to $\Gamma Q_{\theta}$. To increase the learning speed, Vincent et al. (2025b) propose to learn consecutive Bellman iterations in parallel. This approach uses a sequence of online parameters $(\theta_{i})_{i=1}^{K}$ and a sequence of target parameters $(\bar{\theta}_{i})_{i=0}^{K-1}$. Each online network $Q_{\theta_{i+1}}$ is trained to regress $\Gamma Q_{\bar{\theta}_{i}}$. Similarly to DQN, each target parameter $\bar{\theta}_{i}$ is updated to the online parameter $\theta_{i+1}$ every $T$ steps. Importantly, the structure of a chain is enforced by setting each $\bar{\theta}_{i}$ to $\theta_{i}$ every $D\ll T$ steps so that each $Q_{\theta_{i+1}}$, which is learned to regress $\Gamma Q_{\bar{\theta}_{i}}$, are forced to approximate $\Gamma Q_{\theta_{i}}$. This results in $Q_{\theta_{K}}\approx\Gamma Q_{\theta_{K-1}}\approx\ldots\approx\Gamma^{K}Q_{\theta_{0}}$, thus learning $K$ consecutive Bellman iterations in parallel. Importantly, DQN can be recovered by setting $K=1$. While the feature representation can be shared across the online $Q$-networks, iterated $Q$-Network (i-QN) has the drawback of requiring an old copy of the online networks to stabilize training, significantly increasing the memory footprint. In the following, we will explain how the concept of i-QN can help reduce the performance gap between target-free and target-based approaches while maintaining a low memory footprint.
@@ 3 Related Work
Alternatively, some works construct the regression target from the online network instead of the target network, but still use a target network in some other way. For example, Ohnishi et al. (2019) compute the TD(0) loss from the online network and add a term in the loss to constrain the predictions of the online network for the next state-action pair $(s^{\prime},a^{\prime})$ to remain close to the one predicted by the target network. Piché et al. (2021; 2023) develop a similar approach, enforcing similar values for the state-action pair $(s,a)$. Lindström et al. (2025) show that the target network can be removed after a pretraining phase in which they rely on expert demonstrations. Hendawy et al. (2026) construct the target by using the lowest estimate between the online and target networks’ predictions to learn from up-to-date estimates without suffering from the overestimation bias.
@@
Many regularization techniques have been developed, attempting to combat the performance drop that occurs when removing the target network. We stress that our approach is orthogonal to these regularization techniques and we show in Section 5 that our method improves the performance of target-free methods equipped with these advancements. Li and Pathak (2021) encode the input of the $Q$-network with learned Fourier features. While this approach seems promising, the authors acknowledge that the performance degrades for high-dimensional problems. Shao et al. (2022) remove the target-network and search for an action that maximizes the $Q$-network predictions more than the action proposed by the policy. Searching for a better action requires additional resources and is only relevant for actor-critic algorithms. Kim et al. (2019) leverage the MellowMax operator to get rid of the target network. However, the temperature parameter needs to be tuned (Kim, 2020), which increases the compute budget, and a follow-up work demonstrates that the reintegration of the target network is beneficial (Gan et al., 2021). We combine the MellowMax operator with the presented approach in Section D.1 and demonstrate that it bridges the performance gap between the target-free and target-based approaches. Finally, Bhatt et al. (2024) point out the importance of using batch normalization (Ioffe and Szegedy, 2015) to address the distribution shift of the input given to the critic. Our investigation reveals that it degrades the performance in a discrete action setting (see Figure 21, right).
@@
The idea of learning multiple Bellman iterations has been introduced by Schmitt et al. (2022). They demonstrate convergence guarantees in the case of linear function approximation. Then, Vincent et al. (2024) used this approach to learn a recurrent hypernetwork generating a sequence of $Q$-functions where each $Q$-function approximates the Bellman iteration of the previous $Q$-function. Finally, Vincent et al. (2025b) introduced iterated $Q$-Network as a far-sighted version of DQN that learns the $K$ following Bellman iterations in parallel instead of only learning the following one. While promising, those approaches rely on a separate copy of the learnable parameters to stabilise the training process, which increases the memory footprint. In this work, we propose to leverage the potential of iterated $Q$-learning to boost the learning speed of target-free algorithms.
@@ 4 Method
Our goal is to design a new algorithm that improves the learning speed of target-free value-based RL methods without significantly increasing the number of parameters used by the $Q$-networks. To achieve this, we consider a single $Q$-network parameterized with $K+1$ heads. We note $\omega_{k}$ the parameters of the $k^{\text{th}}$ head, $\omega$ the shared parameters, and define $\theta=(\omega,\omega_{0},..,\omega_{K})$ and $\theta_{k}=(\omega,\omega_{k})$. Following Vincent et al. (2025b), for a sample $d=(s,a,r,s^{\prime})$, the training loss is
@@
In the following, we apply iterated Shared Features to several target-based approaches on multiple RL settings, demonstrating that it reduces the gap between target-free and target-based methods. For each algorithm A, we note TB-A as its target-based version, TF-A as its target-free version, and iS-A as the iterated shared approach, where “iS” stands for iterated Shared. For example, for the experiments with the DQN algorithm, we note TF-DQN, the target-free version, and iS-DQN, the iterated Shared version. Importantly, we incorporate the insights provided by Gallici et al. (2025) to use LayerNorm (Ba et al., 2016) for the experiments with discrete action spaces, as we found it beneficial, even for the target-based approach. Similarly, we use BatchNorm (Ioffe and Szegedy, 2015), as suggested by Bhatt et al. (2024), to improve sample-efficiency in continuous action settings, except for the target-based approach, as it degrades performances (see Figure 25, right).
@@ 5.3 Online Continuous Control
We investigate the behavior of iS-QN on the DeepMind Control suite (Tassa et al., 2018), focusing on the hard tasks. We select Soft Actor-Critic (SAC, Haarnoja et al. (2018)) as the base algorithm and adapt the architecture to the one proposed by Lee et al. (2025) (SimbaV2) so that the target-based approach corresponds to the state-of-the-art. This experiment allows us to test iS-QN on different learning dynamics, as the target updates are done with an exponentially moving average instead of a hard update, and the loss for the critic uses a categorical distribution to learn the distribution of the return. Interestingly, Figure 6 (left) shows that only using an old copy of the last layer of the critic to construct the regression target (iS-SAC $K=1$) recovers the performance drop incurred by the target-free approach compared to the target-based approach. Importantly, Lee et al. (2025) design the critic with significantly more parameters than the actor, as commonly done in the actor-critic literature  (Mysore et al., 2021; Mastikhina et al., 2025). This means that iS-SAC K=1 reduces the total number of parameters by $49\%$, see Figure 6 (right). When considering more heads to learn the following Bellman updates, we find it beneficial to give more importance to the first Bellman updates by scaling the future terms in the loss by a discounting factor of $0.25$. We leave the investigation of finding the best way to weight each term in the loss to future work. In Section D.4, we provide a first direction leveraging the concept of meta-gradient reinforcement learning (Xu et al., 2018) to tune learnable coefficients assigned to each term in the loss during training. We note that in this setting, iS-SAC $K=9$ only performs on par with $K=1$. Nonetheless, iS-SAC $K=9$ is still performing better than the best target-free approach, having overlapping confidence intervals with the target-based approach, which serves as the gold standard, as it requires additional parameters.